5/22: RT PC maintenance -> go to safe mode for all VIS connected Dolphin (BS)
- 5/22-23: MZ release from PSL
- 5/20-24: Exchanging cryo valves.
- 6/8: Engineering run
5/22AM~: tuned off cryostat cooler, finesse measurement for Xarm
5/22AM~: tuned off shield cooler, finesse measurement for Xarm
5/2: Cooling start
- 5/29: Cooling done
5/23AM~: tuned off cryostat cooler, finesse measurement for Yarm
5/17~: tuned off ETMY cryostat cooler
- 5/24PM~ inject 100Pa Nitrogen.
- 5/27 evacuation of Nitrogen.
- 6/3-6/14: Installation of TMS for ETMY
- 4/22-5/31: Yarm commissioning
- 5/27-6/30: DRMI commissioning
- 7/1-8/9: DRFPMI commissioning
- 8/12-10/4: DRFPMI ASC
- 10/7-11/15: Post Commissioning
- 11/18-12/13: Noise hunting
- COM: ETMY pitch has still range limit. One of the motor might reach the edge, and one more motor on the same DOF should work if somebody goes to EY and put more current on the driver.
- There was request to CRY people to attend the morning or noon meeting.
- CRY: Exchanging cryo valve during this week.
Report of work
- CRY: ITMX cryostat cooler and shield cooler were stopped yesterday. Outgas pressure was normal.
- CRY: ITMY cryostat cooler was stopped yesterday. Outgas pressure was normal.
- CRY: ETMY Shield cooler was stopped yesterday and there was a big increment of pressure. It means ETMY may have some contamination.
- IOO: MZ was released from the main path of PSL.
- IOO team is still woking on realignment after removing MZ. Unexpected large beam shift by removing BS happened.
- Type Bp: Cleaned aroundf PR2.
- Type Bp: Stepping motor did not work on PR suspensions GAS filter.
- Type B, Bp: Checking for LVDT for PR2 SR2 and BS
- Type B: It was found that loose cable connection on H3 coil for RM of SR2. It was fixed.
- DGS: We added BIO cards to k1lsc0. However it turned out that rebooting k1lsc always fails. Actually it was needed to shutdown to recover. V3 PC seems to have some problem.
- DGS: Maintenance work for adding BIO cards.
- DGS: Camera server maintenance. If you need camera, please use k1cam1.
- CRY: Repairing work for cooler valve will finish today. The last item is for ITMY shield cooler.
- COM: Finesse will be done tonight or tomorrow morning, depending on progress of IOO work.